# Understanding and Embracing Imperfection in Physical Learning Networks

### v1

@@ Introduction
Imperfections in the physical components can affect the bias $B$ in several ways. Bias can arise from enforcing matching between non-identical twin-networks, inaccurately measuring the free and clamped voltage states, and imprecisely implementing the learning rule. As detailed in Appendix C, the latter effect dominates, creating estimated biases of order 2 V/s per edge, in agreement with the measured $B$ values in the previous section (see Appendix E). Device-to-device variations that cause this effect make a priori prediction of the direction of the bias $\mathbf{B}$ infeasible, so the learning rule per edge cannot be directly corrected. However, next, we introduce a modification to the training protocol that drastically reduces the effects of bias.
Taming Imperfection
Figure 4: Schematic example of standard (left, gray) vs overclamping (right, yellow) for a single datapoint. Standard clamping uses $\eta=1$, which results in clamped output $O^{C}=L$ (label). This technique (with any $\eta$) does not result in the free output settling at the label ($O^{F}\rightarrow L$) because the signal ($O^{C}-O^{F}$) decays with the error $\delta=L-O^{F}$, and eventually bias dominates. In contrast, overclamping sets key difference $O^{C}-O^{F}$ to an approximate constant, and reduces train step time $t_{h}$ proportional to the error. This reduces bias with the error signal, and suppresses its effects.
Figure 5: A Classification results for the two methods. Background color is output after training, dots are training data. Both methods use cycle time $\tau=200~{}\mu$s, further details in Appendix B. The data is divided differently in each row, and columns are in order of increasing input variation $\epsilon$. B Classification accuracy for each method as a function of input variation $\epsilon$. Each data point is an average of the 8 experiments of the same task with rotated labels. Error bars are standard error. C Hinge loss as a function of input variation. Averaging same as in B. Log axis breaks to include 0 (black line).
Our modified protocol, overclamping, reduces the impact of bias simply by modifying the amplitude and duration ($t_{h}$) of the clamped boundary condition for the training set. In Coupled Learning (Eq. (3)), the output voltage is nudged proportionally to the signed error $\delta_{i}=L_{i}-O_{i}^{F}$. As the system approaches the solution, $\delta_{i}\rightarrow 0$ so the bias term dominates. To avoid this problem, we nudge in the same direction but by a constant value $\eta L_{0}$:
@@
For a good choice of $\eta L_{0}$, the voltage difference in the learning rule $(V_{F}^{2}-V_{C}^{2})$ is large enough to dominate the bias $B$. To ensure convergence and stability near zero error, we implement an error-dependent step time $t_{h}=r_{0}|\delta|$. This technique is sketched visually alongside standard clamping in Fig. 4. We modify $r_{0}$ during training to keep the period ($\tau$) approximately constant. In theory, these modifications give qualitatively identical results to standard Coupled Learning, but in practice are far less sensitive to the presence of a nonzero bias, see Appendix F for derivation and details.
We demonstrate results of the overclamping protocol for a set of binary classification tasks. We utilize the same method for classification as in Fig. 1, wherein outputs exceeding their labels are considered to have zero error. For overclamping this amounts to $t_{h}=0$ and thus simply skipping these data in training. We train the system to classify eight datapoints arranged in a circle into two classes, as shown for several examples in Fig. 5A. Coupled Learning (standard clamping) learns passably when input values are widely separated, but as the diameter $\epsilon$ of the training data in input space shrinks, performance rapidly degrades, as shown in Fig. 5B and C. In contrast, overclamping not only succeeds (0% classification error), but creates a far sharper boundary (lower hinge loss), possibly because the signal does not vanish at the solution.
Discussion
@@
Second, the effects of noise require more study; in particular, it is important to understand in what contexts noise may be beneficial. Noise on parameter updates has been shown to increase robustness in deep feed-forward digital [26] and analog networks [27]. Similar effects are seen when randomizing inputs (stochastic gradient descent) [28, 29] or randomly dropping connections (dropout) [30]. Noise in edge update timing has been shown to improve solutions in linear self-learning analog networks [17], but the impact of update noise in these systems is unknown. Currently, noise is negligible for our macroscopic electronic components but could become significant in networks with more components that are microfabricated, or that use different resistive elements like memristors. In such systems, noise could arise both in the updates and node voltages. The effects of and interplay between these sources of noise in Contrastive Local Learning Networks is an important avenue for future work.
@@
We have demonstrated for classification tasks that the harmful effects of too much bias can be mitigated using our system-agnostic technique, overclamping, without any modification to the physical components themselves. Our conceptual approach of controlling step size by step time rather than an ever-shrinking error signal can be useful for training a broad range of physical systems, where component variability, thermal fluctuations (noise), and other imperfections can inhibit training, including spintronics [32], memristors [33, 34], and even mechanical systems [35].
@@ Appendix A Error Definitions
where $\Theta$ is the Heaviside step function, zeroing out any data point for which $L_{i}$ has the same sign as $-\delta_{i}$.
@@
Training for Binary Classification
When training for binary classification, (e.g. Fig. 1C), each class ($+$,--) has a single label with $L_{+}=-L_{-}$. The training protocol is applied as normal, except when an output value exceeds its label in the desired direction, $\left(O_{F}-L\right)\times\text{sign}(L)>0$. In this scenario $\eta$ is set to 0, making $O^{C}=O^{F}$ and (in theory) driving no change in the network. This protocol effectively creates a shifted hinge loss, useful for classification problems.
@@ Appendix B Task Data
In all figures, task $\alpha$ is blue and $\beta$ is red. $V_{+}=0.435$ V, $V_{-}=0.018$ V.
Fig 1C-D: Labels for task $\alpha$ are all $L_{+}=63$ mV, and task $\beta$ are $L_{-}=-63$ mV. Inputs are arranged in a square grid in 2D input space [$V_{1}$,$V_{2}$] as shown.
Fig 1E-F: [$V_{1}$, $L$]: $\alpha$ [$V_{+}$, 0.021]V and $\beta$ [$V_{-}$, -0.063]V.
Fig 2B,F: [$V_{1}$, $L$]: $\alpha$ [$V_{+}$, 0.31]V and $\beta$ [$V_{-}$, 0.14]V.
Fig 2C,G: [$V_{1}$, $L$]: $\alpha$ [$V_{+}$, 0.31]V and $\beta$ [$V_{-}$, 0.18]V.
Fig 3: [$V_{1}$, $L$]: $\alpha$ [$V_{-}$, 0.143]V and [$V_{+}$, 0.268]V and $\beta$ [$V_{-}$, 0.156]V.
Fig 5: Labels for task $\alpha$ are all $L_{+}=18$ mV, and task $\beta$ are $L_{-}=-18$ mV. Overclamping uses $L_{0}=V_{+}$, $\eta=0.25$. Standard clamping uses $\eta=1$. Inputs are arranged equally spaced around a ring, centered in 2D input space [$V_{1}$,$V_{2}$] as shown, with radii equally spaced between ϵ=italic-$\epsilon=$83 mV to 417 mV. Data in panels B and C denote the mean and standard error across eight trials, one with every possible linear split of these data locations (corresponding to rotating the dividing line between red and blue points). Two of the eight such divisions are shown in panel A.
@@ Appendix C Learning Rule and Bias
$A_{0}\equiv\frac{1}{V_{0}R_{0}C_{0}}=\frac{1}{(0.33~{}\textrm{V})(100~{}\Omega)(22~{}\mu\textrm{F})}\approx\frac{1.4}{\textrm{V~{}ms}}$ (13)
@@
$A_{0}=\frac{1}{(0.33~{}\textrm{V}(1~{}\textrm{k}\Omega)(2.2~{}\mu\textrm{F})}\approx\frac{1.4}{\textrm{V~{}ms}}$ (14)
@@
In practice, this rate deviates slightly from its predicted value. We measure $A_{0}$ by enforcing and measuring a variety of voltages on all four terminals of a pair of twin self-adjusting resistors, and measuring the resulting change in gate voltage. By comparing predicted (Eq. 3) and actual change, we find that for the experiments of Fig. 2 (Eq. 13), $A_{\exp}\approx 2A_{0}$, and for the remaining experiments (eq. 14), $A_{\exp}\approx A_{0}\text{ to }1.5A_{0}$. The deviation likely comes from using $R_{0}$ out or nearly out of spec for the analog multiplier, as $R_{\min}=1~{}\textrm{k}\Omega$. We use the appropriate value of $A_{\exp}$ to generate our model predictions in Figs 2 and 3. We note that the only effect of this change is to accelerate the dynamics.
The bias $B$ in Eq. 3, however, does alter the dynamics, as we show in the main text. It comes from imperfect aspects of the electronic components, and likely as a result of the effects detailed below. Notably, in each case the bias will be an undefined, likely complicated but deterministic function of the system state. We note that each source of bias listed below scales as $(R_{0}C_{0})^{-1}$, making a change in these components unhelpful in reducing bias. For simplicity we choose $R_{0}=1~{}\textrm{k}\Omega$, $C_{0}=2.2~{}\mu\textrm{F}$ for calculations below, and note that doing so has no effect on the relative size of the bias sources to each other or the learning rate.
Network Discrepancy: One potential source for bias is the discrepancy between free and clamped networks. Measuring in a variety of scenarios with $\eta=0$, we find deviations in node voltage of order 1 mV, with typical voltage drops of order 100 mV. To probe the maximum effect of this bias we assume the system is in a state where the learning rule would be zero in a perfectly mirrored system (signal = 0). However, discrepancies create deviations such that $V_{F}\approx V_{C}+1$ mV. Plugging this into (3) we have:
[…]
@@
Voltage Measurements: A second potential source is imperfect voltage measurements on-edge, which comes down to component error of both operational amplifiers and a multiplier circuit. These imperfections would generate error in the same manner as the previous calculation, and are of the same order (1% error) based on the specifications of the integrated circuits used (TLV274IN and AD633ANZ).
Imperfect Learning Execution: The third and likely dominant source of bias is the output offset voltage of the multiplier chip (AD633ANZ), which generates the charging current. This bias voltage (listed as $V_{B}\approx$ 5 mV) directly creates a current $V_{B}/R_{0}\approx 5~{}\mu\textrm{A}$ onto the capacitor. Then we have:
[…]
@@ Appendix D Theory
where $\gamma=A_{0}/S$ is the learning rate. As described in Appendix C, $A_{0}$ deviates from its predicted value depending on a choice of electronic components. As a result we measure this effect (see Appendix C) and use the resulting $\gamma\approx 2.4\times 10^{6}~{}\Omega/\textrm{s}$ for the theory results of Fig. 2, and $\gamma=3.2\times 10^{6}~{}\Omega/\textrm{s}$ for Fig. 3.
Second, we linearize the learning dynamics around a minimum $\mathbf{G}^{*}$ of the contrastive function, i.e. a solution.
Here, $\mathbf{M}=\gamma\frac{\partial^{2}\mathcal{C}}{\partial\mathbf{G}\partial\mathbf{G}}\Big{|}_{\mathbf{G}^{*}}$, is proportional to the Hessian of the contrastive function evaluated at the solution.
Third, we consider linear resistor networks: the voltage response minimizes the power dissipation $\mathcal{P}=\sum_{i}K_{i}(\Delta V)_{i}^{2}$ where $K_{i}(G_{i})$ are the conductances as a function of gate voltages, see Eq. (2). This is approximately correct and equivalent to the analytical treatment given to the same system in [20]. As a result, the contrastive function becomes an explicit function of the gate voltages,
@@
from which we can directly compute the contrastive Hessian $\mathbf{M}$. For more details see [25].
@@
[…]
where $\mathbf{G}_{0}$ is the initial gate voltage state and $e^{-\mathbf{M}t}$ denotes the matrix exponential of $-\mathbf{M}t$. Considering that $\mathbf{M}$ is a positive semi-definite matrix, having only positive or zero eigenvalues, we can further decompose the previous relation in terms of its (pseudo)inverse $\mathbf{M}^{\dagger}$ and projector onto its nullspace $\mathbf{N}$
$\mathbf{G}(t,\mathbf{G}_{0})=e^{-\mathbf{M}t}\mathbf{G}_{0}+(1-e^{-\mathbf{M}t})(\mathbf{G}^{*}+\mathbf{M}^{\dagger}\mathbf{B})+t\mathbf{N}\mathbf{B}.$ (21)
@@
Periodic training: Equation (21) remains valid for periodic training, where at each half-period both the Hessian and the solution change with the task, $M_{\mu}$ and $\mathbf{G}^{*}_{\mu}$, with $\mu=\{\alpha,\beta\}$. Denoting the evolution of the gate voltages under each task by $\mathbf{G}_{\mu}(t)$, the analytical solution under periodic training is found by imposing continuity in time,
@@
$\bar{\mathbf{G}}_{\text{0}}=\left(\mathbf{I}-\mathbf{A}_{\beta}\mathbf{A}_{\alpha}\right)^{-1}\left(\mathbf{A}_{\beta}\mathbf{b}_{\alpha}+\mathbf{b}_{\beta}\right),$ (27)
with $\mathbf{I}$ the identity matrix. The inverse exists whenever there are closed orbits (e.g. Figs. 2 and  3), otherwise it should be taken as a pseudoinverse. Finally, the evolution of the gate voltages over a full period can be found using $\bar{\mathbf{G}}_{0}$ as the initial condition into the dynamical equation 21 for task $\alpha$, chained with those for task $\beta$.

### v2

@@ II.4 Matching Experiment and Theory
Imperfections in the physical components can affect the bias $B$ in several ways. Bias can arise from enforcing matching between non-identical twin-networks, inaccurately measuring the free and clamped voltage states, and imprecisely implementing the learning rule. As detailed in Materials and Methods, the latter effect dominates, creating estimated biases of order 2 V/s per edge, in agreement with the measured $B$ values in the previous section. Device-to-device variations that cause this effect make a priori prediction of the direction of the bias $\mathbf{B}$ infeasible, so the learning rule per edge cannot be directly corrected. However, next, we introduce a modification to the training protocol that drastically reduces the effects of bias.
II.5 Taming Imperfection
Figure 4: Standard vs Overclamping Schematic example of standard (left, gray) vs overclamping (right, yellow) for a single datapoint. Standard clamping uses $\eta=1$, which results in clamped output $O^{C}=L$ (label). This technique (with any $\eta$) does not result in the free output settling at the label ($O^{F}\rightarrow L$) because the signal ($O^{C}-O^{F}$) decays with the error $\delta=L-O^{F}$, and eventually bias dominates. In contrast, overclamping sets the key difference $O^{C}-O^{F}$ to an approximate constant, and reduces train step time $t_{h}$ proportional to the error. This reduces bias with the error signal, and suppresses its effects.
[…]
Our modified protocol, overclamping, reduces the impact of bias simply by modifying the amplitude and duration ($t_{h}$) of the clamped boundary condition for the training set. In Coupled Learning (Eq. (3)), the output voltage is nudged proportionally to the signed error $\delta_{i}=L_{i}-O_{i}^{F}$. As the system approaches the solution, $\delta_{i}\rightarrow 0$ so the bias term dominates. This has the effect of shifting the learning landscape (resulting in an error floor) and pushing the system along basins of zero error, which creates the demonstrated cyclic dynamics. To avoid these problems, we nudge in the same direction but by a constant value $\eta L_{0}$:
@@ II.5 Taming Imperfection
For a good choice of $\eta L_{0}$, the voltage difference in the learning rule $(V_{F}^{2}-V_{C}^{2})$ is large enough to dominate the bias $B$. To ensure convergence and stability near zero error, we implement an error-dependent step time $t_{h}=t_{0}|\delta|$. This technique is sketched visually alongside standard clamping in Fig. 4. We modify $t_{0}$ during training to keep the period ($\tau$) approximately constant. In theory, these modifications give qualitatively identical results to standard Coupled Learning, but in practice are far less sensitive to the presence of a nonzero bias, see Materials and Methods for derivation and details.
We demonstrate results of the overclamping protocol for a set of binary classification tasks. We utilize the same method for classification as in Fig. 1, wherein outputs exceeding their labels are considered to have zero error. For overclamping this amounts to $t_{h}=0$ and thus simply skipping these data in training. With this different relationship to training time, a direct comparison of the steady-state results of Fig. 1 is difficult, however analysis of the final results is directly comparable. We note also that overclamping shines with smaller labels and small input variation, whereas standard clamping requires both to be larger to even sense the error signal. To highlight a general comparison, we train the system to classify eight datapoints arranged in a circle (with trials for five different radii) into two classes, as shown for several examples in Fig. 5A. Coupled Learning (standard clamping) learns passably when input values are widely separated, but as the diameter $\epsilon$ of the training data in input space shrinks, performance rapidly degrades, as shown in Fig. 5B and C. In contrast, overclamping not only succeeds (0% classification error), but creates a far sharper boundary (lower hinge loss), possibly because the signal does not vanish at the solution.
III Discussion
@@
Second, the effects of noise require more study; in particular, it is important to understand in what contexts noise may be beneficial. Noise on parameter updates has been shown to increase robustness in deep feed-forward digital [28] and analog networks [29]. Similar effects are seen when randomizing inputs (stochastic gradient descent) [30, 31] or randomly dropping connections (dropout) [32]. Noise in edge update timing has been shown to improve solutions in linear self-learning analog networks [18], but the impact of update noise in these systems is unknown. Currently, noise is negligible for our macroscopic electronic components but could become crucial in networks with more components that are microfabricated, or that use different resistive elements like memristors. In such systems, noise could arise both in the updates and node voltages. The effects of and interplay between these sources of noise in Contrastive Local Learning Networks is an important avenue for future work. Importantly, overclamping also can be used to suppress the effects of noise by amplifying the learning signal.
@@
We have demonstrated for classification tasks that the harmful effects of too much bias can be mitigated using our system-agnostic technique, overclamping, without any modification to the physical components themselves. Our conceptual approach of controlling step size by step time rather than an ever-shrinking error signal can be useful for training a broad range of physical systems, where component variability, thermal fluctuations (noise), and other imperfections can inhibit training, including spintronics [34], memristors [35, 36], and even mechanical systems [37]. Notably, in analog systems, the steps of measurement, calculation and application of updates are all potentially noise- and bias-laden, and each benefit from enhanced signals. Furthermore, it is likely more energy-efficient than standard Coupled Learning, since the change in power from the modified output boundary conditions is small and varied in sign, but the time each state is run is nearly always reduced. Moreover, overclamping is a broadly applicable technique that extends beyond Coupled Learning. By modifying the duration and magnitude of the boundary conditions, it can help counterbalance imperfect learning signals that hinder (physical) learning rules, such as directed aging [38], equilibrium propagation [20] and its variants, and any training protocol that uses perturbations to extract gradient information, including of weights, as in direct feedback alignment or multiplexed gradient descent [39].
@@ Acknowledgements
Data Availability
All data and software are publicly available at https://doi.org/10.5281/zenodo.15692914 [46].
@@ Appendix A Error Definitions
where $\Theta$ is the Heaviside step function, zeroing out any data point for which $L_{i}$ has the same sign as $\delta_{i}$.
@@
When training for binary classification, (e.g. Fig. 1C), each class (++,-) has a single label with $L_{+}=-L_{-}$. The training protocol is applied as normal, except when an output value exceeds its label in the desired direction, $\left(O_{F}-L\right)\times\text{sign}(L)>0$. In this scenario $\eta$ is set to 0, making $O^{C}=O^{F}$ and (in theory) driving no change in the network. This protocol effectively creates a shifted hinge loss, useful for classification problems.
@@ Appendix B Task Data
In all figures, task $\alpha$ is blue and $\beta$ is red. $_{+}=0.435 V, V$$_{-}=0.018 V$.
Fig 1C-D: Labels for task $\alpha$ are all $L_{+}=63$ mV, and task $\beta$ are $L_{-}=-63$ mV. Inputs are arranged in a square grid in 2D input space $V_{1}$$V_{2}$] as shown.
Fig 1E-F: $V_{1}$, $L$]: $\alpha$ $V_{+}$, 0.021]V and $\beta$ $V_{-}$, -0.063]V.
Fig 2B,F: $V_{1}$, $L$]: $\alpha$ $V_{+}$, 0.31]V and $\beta$ $V_{-}$, 0.14]V.
Fig 2C,G: $V_{1}$, $L$]: $\alpha$ $V_{+}$, 0.31]V and $\beta$ $V_{-}$, 0.18]V.
Fig 3: $V_{1}$, $L$]: $\alpha$ $V_{-}$, 0.143]V and $V_{+}$, 0.268]V and $\beta$ $V_{-}$, 0.156]V.
Fig 5: Labels for task $\alpha$ are all $L_{+}=18$ mV, and task $\beta$ are $L_{-}=-18$ mV. Overclamping uses $L_{0}=V_{+}$, $\eta=0.25$. Standard clamping uses $\eta=1$. Inputs are arranged equally spaced around a ring, centered in 2D input space $V_{1}$$V_{2}$] as shown, with radii equally spaced between $\epsilon=$83 mV to 417 mV. Data in panels B and C denote the mean and standard error across eight trials, one with every possible linear split of these data locations (corresponding to rotating the dividing line between red and blue points). Two of the eight such divisions are shown in panel A.
@@ Appendix C Learning Rule and Bias
$A_{0}\equiv\frac{1}{V_{0}R_{0}C_{0}}=\frac{1}{(0.33~\textrm{V})(100~\text{ohms})(22~\mu\textrm{F})}\approx\frac{1.4}{\textrm{V~ms}}$ (13)
@@
$A_{0}=\frac{1}{(0.33~\textrm{V}(1~\textrm{k~ohms})(2.2~\mu\textrm{F})}\approx\frac{1.4}{\textrm{V~ms}}$ (14)
@@
In practice, this rate deviates slightly from its predicted value. We measure $A_{0}$ by enforcing and measuring a variety of voltages on all four terminals of a pair of twin self-adjusting resistors, and measuring the resulting change in gate voltage. By comparing predicted (Eq. 3) and actual change, we find that for the experiments of Fig. 2 (Eq. 13), $A_{\exp}\approx 2A_{0}$, and for the remaining experiments (eq. 14), $A_{\exp}\approx A_{0}\text{ to }1.5A_{0}$. The deviation likely comes from using $R_{0}$ out or nearly out of spec for the analog multiplier, as $R_{\min}=1~\textrm{k~ohms}$. We use the appropriate value of $A_{\exp}$ to generate our model predictions in Figs 2 and 3. We note that the only effect of this change is to accelerate the dynamics.
The bias $B$ in Eq. 3, however, does alter the dynamics, as we show in the main text. It originates from imperfect aspects of the electronic components, and likely as a result of the effects detailed below. For each edge the bias is an unknown, likely complicated but deterministic function of the system state. We note that each source of bias listed below scales as $(R_{0}C_{0})^{-1}$, making a change in these components unhelpful in reducing bias. For simplicity we choose $R_{0}=1~\textrm{k~ohms}$, $C_{0}=2.2~\mu\textrm{F}$ for calculations below, and note that doing so has no effect on the relative size of the bias sources to each other or the learning rate.
Non-identical Twin Networks: One potential source of bias is the discrepancy between free and clamped networks. For a variety of voltages applied at inputs, in the absence of clamping at the outputs, the free and clamped networks should have the same set of node voltages. We find deviations in node voltages of order 1 mV, with typical voltage drops of order 100 mV. To probe the maximum effect of this bias we assume the system is in a state where the learning rule would be zero in a perfectly mirrored system (signal = 0). However, discrepancies create deviations such that $V_{F}\approx V_{C}+1$ mV. Plugging this into (3) we have:
$\Delta G=t_{h}\times A_{0}\Big(201~(\textrm{mV})^{2}\Big)=t_{h}\times(275~\textrm{mV/s}),$ (15)
@@
Imperfect Voltage Drop Measurements: A second potential source of bias is imperfect voltage drop measurements across each edge, which comes down to component error of operational amplifiers and a multiplier circuit. These imperfections would generate error in the same manner as the previous calculation, and are of the same order (1% error) based on the specifications of the integrated circuits used (TLV274IN and AD633ANZ).
Imperfect Execution of Learning Rule: The third and likely dominant source of bias is the output offset voltage of the multiplier chip (AD633ANZ), which generates the charging current. This bias voltage (listed as $V_{B}\approx$ 5 mV) directly creates a current $V_{B}/R_{0}\approx 5~\mu\textrm{A}$ onto the capacitor. Then we have:
$\Delta G=t_{h}\times\frac{V_{B}}{R_{0}C_{0}}=t_{h}\times(2.3~\textrm{V/s})$ (16)
@@ Appendix D Theory
where $\gamma=A_{0}/S$ is the learning rate. As described in the previous section, $A_{0}$ deviates from its predicted value depending on a choice of electronic components. As a result we measure this effect and use the resulting $\gamma\approx 2.4\times 10^{6}~\text{ohm}/\textrm{s}$ for the theory results of Fig. 2, and $\gamma=3.2\times 10^{6}~\text{ohm}/\textrm{s}$ for Fig. 3.
Second, we linearize the learning dynamics around a minimum $\mathbf{G}^{*}$ of the contrastive function, i.e. a learning solution,
$\frac{d\mathbf{G}}{dt}=-\mathbf{M}(\mathbf{G}-\mathbf{G}^{*})+\mathbf{B}.$ (18)
Here, $\mathbf{M}=\gamma\frac{\partial^{2}\mathcal{C}}{\partial\mathbf{G}\partial\mathbf{G}}\Big|_{\mathbf{G}^{*}}$, is proportional to the Hessian of the contrastive function evaluated at the solution. The choice of $\mathbf{G}^{*}$ is determined by the next approximation.
Third, we linearize the physical dynamics around the zero voltage state—for this calculation we emphasize that we consider linear resistor networks. We recall that the conductance of a single edge $i$ (Eq. (2)) connecting nodes $a$ and $b$ depends on the voltage at both terminals, $K_{i}=S(G_{i}-V_{T}-0.5(V_{a}+V_{b}))$, where $S$ and $V_{T}$ are fixed values. The power dissipated by this edge, expanded to quadratic order in voltages, corresponds to $K_{i}(V_{a}-V_{b})^{2}\approx S(G_{i}-V_{T})(V_{a}-V_{b})^{2}$. In other words, in the linear regime conductances become independent of the voltages $V_{a}$ and $V_{b}$. Moreover, the voltage response to some fixed voltage inputs is defined by Kirchhoff’s current law, and corresponds to the minimization of the power dissipation $\mathcal{P}=\sum_{i}S(G_{i}-V_{T})(\Delta V)_{i}^{2}$. This is approximately correct and equivalent to the analytical treatment given to the same system in [21]. Following [27], the linearized physical response to external current is simply
$\mathbf{V}_{F}=\mathbf{H}^{-1}\mathbf{I},$ (19)
[…]
As a result, the contrastive function becomes an explicit yet lengthy function of the gate voltages,
@@
from which we can directly compute the contrastive Hessian $\mathbf{M}$. For more details see [27] and the code repository [46].
In practice, we choose $\mathbf{G}^{*}$ to be the point in the solution space of the linear resistor network that lies closest to the location in gate voltage space where the bias is measured (see next section).
@@
$\mathbf{G}(t,\mathbf{G}_{0})=e^{-\mathbf{M}t}\mathbf{G}_{0}+\left[\int_{0}^{t}e^{-\mathbf{M}s}ds\right](\mathbf{M}\mathbf{G}^{*}+\mathbf{B}),$ (21)
where $\mathbf{G}_{0}$ is the initial gate voltage state and $e^{-\mathbf{M}t}$ denotes the matrix exponential of $\mathbf{M}t$. Considering that $\mathbf{M}$ is a positive semi-definite matrix, having only positive or zero eigenvalues, we can further decompose the previous relation in terms of its (pseudo)inverse $\mathbf{M}^{\dagger}$ and projector onto its nullspace $\mathbf{N}$
$\mathbf{G}(t,\mathbf{G}_{0})=e^{-\mathbf{M}t}\mathbf{G}_{0}+(\mathbf{1}-e^{-\mathbf{M}t})(\mathbf{G}^{*}+\mathbf{M}^{\dagger}\mathbf{B})+t\mathbf{N}\mathbf{B}.$ (22)
@@
Periodic training: Equation (22) remains valid for periodic training that alternates between tasks $\alpha$ and $\beta$. At each half-period both the Hessian and the solution change with the task, $M_{\mu}$ and $\mathbf{G}^{*}_{\mu}$, with $\mu=\{\alpha,\beta\}$. Denoting the evolution of the gate voltages under each task by $\mathbf{G}_{\mu}(t)$, the analytical solution under periodic training is found by imposing continuity in time,
@@
In Fig. 6 we compare the transient dynamics of the biased and ideal systems modeling the experimental setup of Fig. 2F for three values of $\tau$. Starting from the same initial condition, the learning trajectories in the absence of bias (ideal) reach the joint solution by zigzagging between the solution lines of each task. In contrast, the biased trajectories cannot approach the joint solution and instead get trapped in limit cycles.
Figure 6: Transient dynamics of biased and ideal systems. Learning trajectories in the linear approximation for a biased (black) and ideal (yellow) systems, for three different training periods: $\tau=20\text{ms}$ (left), $80\text{ms}$ (middle), and $200\text{ms}$ (right). The parameters of the network, together with the measured biased (black arrow), are the same as those of Fig. 2F. For each case, we use the same initial condition $(G_{1},G_{2})=(4\text{V},4\text{V})$ and computed the evolution for 20 learning cycles, i.e. a total time of $T=20\tau$.
@@
$\bar{\mathbf{G}}_{\text{0}}=\left(\mathbf{1}-\mathbf{A}_{\beta}\mathbf{A}_{\alpha}\right)^{-1}\left(\mathbf{A}_{\beta}\mathbf{b}_{\alpha}+\mathbf{b}_{\beta}\right).$ (28)
The inverse exists whenever there are closed orbits (e.g. Figs. 2 and  3), otherwise it should be taken as a pseudoinverse. Finally, the evolution of the gate voltages over a full period can be found using $\bar{\mathbf{G}}_{0}$ as the initial condition into the dynamical equation 22 for task $\alpha$, chained with those for task $\beta$.
Analytical solution for the small network
For the three edge network analyzed in the main text (Fig. 2A), we can explicitly write down the inverse of the Physical Hessian in terms of the gate voltages and solve for the output voltage according to Eq. (19). Alternatively, for such a small system we can directly solve by imposing current conservation. Using the notation of the main text (see Fig. 2A), we have $\mathbf{V}_{F}=(V_{+},V_{-},V_{1},O)$, where the first two entries are the high and low ground voltages, the third is the input voltage, and $O$ is the voltage at the output node given by
$O=\frac{V_{1}R_{0}^{-1}+S(G_{2}V_{+}+G_{1}V_{-}-(V_{-}+V_{+})V_{T})}{R_{0}^{-1}+S(G_{2}+G_{1}-2V_{T})},$ (29)
where $G_{1}$ is the gate voltage of the transistor connecting $O$ to $V_{-}$, $G_{2}$ from $O$ to $V_{+}$, and $R_{0}$ is the fixed resistance of the edge connecting $V_{1}$ to $O$.